# Safe Reinforcement Learning on the Constraint Manifold: Theory and Applications

### v1

@@ V-C Drift Clipping
As described in (12), the first term $-{\bm{J}}_{u}^{\dagger}{\bm{\psi}}$ compensates for the constraint drift caused by the system drift. Let’s consider the effect of the system drift when no controller is applied, i.e., $[{\bm{u}}_{s}\;{\bm{u}}_{\mu}]^{\intercal}={\bm{0}}$.
$\dot{\bm{c}}_{i}$ $=\dot{\bm{k}}_{i}+\dot{\bm{\mu}}_{i}={\bm{J}}_{k,i}\bm{f}+{\bm{J}}_{k,i}\bm{G}{\bm{u}}_{s}+\bm{A}_{i}{\bm{u}}_{\mu}$
$={\bm{\psi}}_{i},\quad\quad i\in\{1,\dots,K\}$
Starting from a safe state where $\bm{k}<{\bm{0}}$, we can obtain that when $\psi_{i}>0$, we have $\dot{k}_{i}>0$, indicating that the drift term is pushing the state to the boundary of the constraint $\partial\mathcal{M}$. On the other hand, when $\psi_{i}<0$, the drift term pulls the state away from the constraint boundary. Therefore, the safe controller only needs to compensate for the drift that tends to the boundary. This can be done by clipping the drift term to be non-negative.
@@
We will illustrate the benefit of drift clipping in Example 7 after introducing the extension of separable state space (Section VI-A).
@@ VI Extensions
We have introduced the fundamental concept of ATACOM in Section IV. In this section, we will extend the application scope of the approach in several directions. First, in the previous section, we assumed that we had full knowledge of dynamic systems and that the constraints were determined purely by the system’s state variables. However, in many applications, the constraints are determined by both the state of the robot and the external state, yet we only partially know the dynamic system. For example, the safety constraints in HRI can be defined by the distance between the human and the robot, but the model of the human motion is difficult to obtain. Section VI-A will discuss extending ATACOM for scenarios with only partial knowledge of the dynamics. Next, we introduce an extension to the second-order system in Section VI-B. We show by an example why ATACOM controller does not guarantee safety for a second-order system with position-based constraint and then provide a simple solution that modifies the constraint to ensure safety. This method can also be extended to higher-order systems. Section VI-C discusses how ATACOM can be applied to the problem with equality constraints. The summary of the extensions is shown in Table VI.
@@
Sec. Order $\begin{bmatrix}\bm{s}\\ \dot{\bm{s}}\end{bmatrix}$ $\begin{matrix}\beta(k(\bm{s}))+J_{k}(\bm{s})\dot{\bm{s}}\\ +\bm{\mu}\end{matrix}$ $\begin{bmatrix}\dot{\bm{s}}\\ f(\bm{s},\dot{\bm{s}})\end{bmatrix}+\begin{bmatrix}{\bm{0}}\\ G(\bm{s},\dot{\bm{s}})\end{bmatrix}{\bm{u}}_{s}$ $[{\bm{J}}_{k}\bm{G}\;\;\bm{A}]$ $\begin{matrix}{\bm{J}}_{k}\bm{f}+({\bm{J}}_{\beta}{\bm{J}}_{k}\\ +\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}\end{matrix}$
@@ VI-A Dynamical Environment with Separable State Space
Similar to (11), the following requirement should be satisfied to ensure safety
@@
Compared to the previous derivation (11), the drift term $\psi(\bm{q},\bm{z})$ has an additional source from the uncontrollable state’s motion $J_{z}(\bm{q},\bm{z})\dot{\bm{z}}$. We assume $\dot{\bm{z}}$ can be observed or estimated. In practice, we can estimate the velocity of the uncontrollable states by the finite difference or by using a state observer. A comparison of different velocity observations is shown in Section VII-B. We can derive the ATACOM-controller similarly to Eq. (12) where ${\bm{J}}_{u}$ and ${\bm{\psi}}$ should be adapted. Here, we show a simple example with a dynamic moving obstacle in the 2D environment, illustrating the effect of drift clipping.
@@ VI-B Second-Order Dynamics
Previous derivation focuses on the first-order affine control system with state constraints. However, this method can not be applied directly to higher-order systems by simply converting the system to the first order. We first illustrate from an example that the constraint with only position input violates the assumption. Then, we provide a solution for the second-order systems. This approach can be easily extended to higher-order systems.
@@
Consider a system directly controlling acceleration, i.e., $\ddot{\bm{s}}={\bm{u}}_{s}$. We convert the system to a first-order affine system as
@@
For the constraint $k(\bm{s})<{\bm{0}}$ that only takes $\bm{s}$ as input, the Jacobian with respect to $\widehat{\bm{s}}$ is ${\bm{J}}_{k}=\begin{bmatrix}\frac{\partial}{\partial s}k&{\bm{0}}\end{bmatrix}$. We can verify that $\mathrm{rank}({\bm{J}}_{G})=\mathrm{rank}({\bm{J}}_{k}\bm{G})=0$ where the Assumption 4 does not hold. ∎
Generally speaking, the second-order control affine system can be defined as
@@
$k^{*}(\bm{s},\dot{\bm{s}})=\beta(k(\bm{s}))+\dot{k}(\bm{s})=\beta(k(\bm{s}))+{\bm{J}}_{k}\dot{\bm{s}}\leq{\bm{0}}$ (17)
where $\beta(\cdot)$ is a class $\mathcal{K}$ function, following the idea of High-order Control Barrier Function (CBF) from [35]. We again introduce the slack variable $\bm{\mu}$ to construct the constraint manifold and set the time-derivatives of the new constraint $\bm{k}^{*}$ to zero and obtain
$({\bm{J}}_{\beta}+\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}+{\bm{J}}_{k}\bm{f}+{\bm{J}}_{k}\bm{G}{\bm{u}}_{s}+\bm{A}\bm{\mu}={\bm{0}}$
Grouping the terms together, we obtain the same form of requirement as in Eq. (11). The Jacobian ${\bm{J}}_{u}$ and the drift ${\bm{\psi}}$ are defined as
${\bm{J}}_{u}=[{\bm{J}}_{k}\bm{G}\;\;\bm{A}],\;{\bm{\psi}}={\bm{J}}_{k}\bm{f}+({\bm{J}}_{\beta}{\bm{J}}_{k}+\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}$
@@ VI-C Equality Constraints
The ATACOM controller can be extended for equality constraint
$l(\bm{s})={\bm{0}},$
where $l:\mathcal{S}\rightarrow\mathbb{R}^{L}$ is of class $C^{1}$ with the Jacobian $J_{l}:\mathcal{S}\rightarrow\mathbb{R}^{L\times S}$. We assume the system with equality constraint is under-constrained, i.e., $L<S$. We can formulate the constraint manifold
@@
We can again construct the ATACOM controller using Eq. (12). The controlled system will approach the constraint manifold $\mathcal{M}$ from the neighborhood $\Omega_{\eta}$ and stay on the manifold.
However, simply choosing a matrix ${\bm{T}}\in\mathbb{R}^{N\times U}$ whose diagonal entry is ${\bm{T}}_{[i,i]}=1$ is no longer a valid choice for the coordinate space in this setting. For example, for a 1-sphere in 2D, the coordinate frame ${\bm{T}}=\begin{bmatrix}1&0\end{bmatrix}^{\intercal}$ is not a valid coordinate as there are multiple points mapped to the same point. The 2-sphere does not even contain a smooth varying basis, according to the Hairy Ball Theorem [74]. Unfortunately, the choice of the reference matrix is not trivial and depends on the manifold. We will leave this discussion to future work.
@@ VII Experiments
As shown in the previous works, ATACOM has been applied to solve various tasks with different environmental characteristics, such as Robot Air Hockey that contains equality constraints, collision avoidance with dynamically moving obstacles [9], collision avoidance with differential drive (control affine system), and Human-Robot Interaction (high-dimensional) [10], and it outperforms state-of-the-art baselines.
In this section, we will analyze the effectiveness of different techniques introduced in previous sections in some simple environments to provide a thorough understanding of the algorithm. In the first experiment, we will compare the effect of different slack dynamics functions and their hyperparameters in a 2D static collision avoidance environment (Fig. 8a). Then, we will compare the effects of different velocity observation methods in a dynamic environment containing moving obstacles (Fig. 8b). Additionally, we evaluate how the dynamic mismatch will impact the performance of ATACOM in a simulated air hockey task (Fig. 8c). Finally, we show in the real-world robot air hockey experiment (Fig. 8d) that we can fine-tune the RL agent from online real-world interactions with the help of ATACOM.
@@
VII-A Comparison of Different Slack Dynamics
We introduced multiple types of slack dynamics functions to ensure safety in Section V-A. In this experiment, we will compare the effect of different types of slack dynamics functions (Linear and Exponential) and their hyperparameters $\beta$ in a 2D collision avoidance environment, shown in Figure 8a. A circular obstacle shown in red stays in the middle of the environment. The task is to control the planar robot (blue point) to reach the target (green point). The robot is controlled by velocity
$\begin{bmatrix}\dot{x}\\ \dot{y}\end{bmatrix}=\begin{bmatrix}u_{x}\\ u_{y}\end{bmatrix}$
The constraints are collision avoidance w.r.t. the fixed obstacle, i.e., $\|{\bm{p}}_{r}-{\bm{p}}_{o}\|>r_{\text{o}}$, and keeping the robot inside the boundaries, i.e., $l_{i}<p_{r,i}<u_{i},i\in\{x,y\}$. The reward is the negative distance to the target $r=-\|{\bm{p}}_{r}-{\bm{p}}_{t}\|$. We use SAC [75] as the RL algorithm. The control frequency is set to be $100Hz$ in all experiments. The learning parameters are the same among all experiments except for the parameters in comparison, i.e., the type of slack dynamics function and its parameters $\beta$. Details of the hyperparameters can be found in Appendix A-A. We ran 25 seeds for each experiment setting.
@@ VII-A Comparison of Different Slack Dynamics
The learning curves are shown in Figure 9. In all experiment settings, the robot has not collided with the obstacle during training. The robot can reach the target in all experiment settings within a fixed horizon length except for the setting of linear slack dynamic function with $\beta=0.3$. The reason is that the action space is heavily morphed by the slack dynamics function and the robot is taking a conservative behavior. The lower reward result indicates that the robot is reaching the target slower than the higher reward result. We can clearly see that the slack dynamics function with less morphing of the action space (higher $\beta$ and exponential slack dynamics function) leads to better training performance. However, a higher $\beta$ parameter requires a higher control frequency as the manifold will have a higher curvature, causing the system to deviate from the manifold due to the discretization error.
@@
VII-B Comparison of Different Velocity Observations in Dynamical Environment
@@
We use a 2D Dynamic Environment as a study example, as shown in Fig. 8b. The robot (blue point) is controlled by velocity. The constraints are defined as keeping the distance to each obstacle bigger than a threshold, i.e., $\|{\bm{p}}_{r}-{\bm{p}}_{o_{i}}\|>r_{o},i=1,2,\cdots$. The target (green square) is randomly placed in the environment for each episode. We consider two types of moving obstacles: (1) The obstacles move in a fixed pattern, i.e., each obstacle moves along a circle around its initial position. (2) The obstacles move randomly in the environment. In this experiment, we use a hand-crafted policy that tries to reach the target with a linear tracker
@@
In this experiment, we compare the safety performance of (a) exact velocity observations (EXACT), (b) no zero velocity (NONE), and (c) velocities obtained from the finite difference in position (FD). In the FD setting, we added a Gaussian noise with a standard deviation of $0.03$ to the position. We evaluate the performance with different numbers of obstacles (2, 6, 10) and different obstacle speeds. In the LOW-velocity setting, the maximum speed of the obstacles is $50\%$ of the maximum speed of the robot. The MEDIUM and HIGH-velocity settings are $100\%$ and $150\%$, respectively. We ran 1000 episodes for each experiment setting. The experiment hyperparameters can be found in Appendix A-B.
@@
VII-C Dynamic Mismatch in Air Hockey Simulation

### v2

@@ V-C Drift Clipping
As described in (12), the first term $-{\bm{J}}_{u}^{\dagger}{\bm{\psi}}$ compensates for the constraint drift caused by the system drift. Let’s consider the effect of the system drift when no control is applied, i.e., $[{\bm{u}}_{s}\;{\bm{u}}_{\mu}]^{\intercal}={\bm{0}}$.
$\dot{\bm{c}}_{i}$ $=\dot{\bm{k}}_{i}+\dot{\bm{\mu}}_{i}={\bm{J}}_{k,i}\bm{f}+{\bm{J}}_{k,i}\bm{G}{\bm{u}}_{s}+\bm{A}_{i}{\bm{u}}_{\mu}={\bm{\psi}}_{i},$
for $i\in\{1,\cdots,K\}$. Starting from a safe state where $\bm{k}<{\bm{0}}$, we can obtain that when $\psi_{i}>0$, we have $\dot{k}_{i}>0$, indicating that the drift term is pushing the state to the boundary of the constraint $\partial\mathcal{M}$. On the other hand, when $\psi_{i}<0$, the drift term pulls the state away from the constraint boundary. Therefore, the safe controller only needs to compensate for the drift that tends to the boundary. This can be done by clipping the drift term to be non-negative.
@@
We illustrate the benefit of drift clipping in Example 7 after the extension of separable state space (Section VI-A).
@@ VI Extensions
We introduced the fundamental concept of ATACOM in Section IV. In this section, we will expand the scope of its application in several directions. Previously, we assumed full knowledge of the dynamic system and that the constraints were solely determined by the system’s state variables. However, in many real-world applications, constraints are influenced by both the robot’s state and external states, and we often have only partial knowledge of the dynamic system. For example, the safety constraints in HRI can be defined by the distance between the human and the robot, but the model of the human motion is difficult to obtain. Section VI-A will discuss extending ATACOM for scenarios with only partial knowledge of the dynamics. Next, we introduce an extension to the second-order system in Section VI-B. We show by an example why ATACOM controller does not guarantee safety for a second-order system with position-based constraint and then provide a simple solution that modifies the constraint to ensure safety. Section VI-C discusses how ATACOM can be applied to the problem with equality constraints. The summary of the extensions is shown in Table VI.
@@
Sec. Order $\begin{bmatrix}\bm{s}\\ \dot{\bm{s}}\end{bmatrix}$ $\begin{matrix}\zeta(k(\bm{s}))+J_{k}(\bm{s})\dot{\bm{s}}\\ +\bm{\mu}\end{matrix}$ $\begin{bmatrix}\dot{\bm{s}}\\ f(\bm{s},\dot{\bm{s}})\end{bmatrix}+\begin{bmatrix}{\bm{0}}\\ G(\bm{s},\dot{\bm{s}})\end{bmatrix}{\bm{u}}_{s}$ $[{\bm{J}}_{k}\bm{G}\;\;\bm{A}]$ $\begin{matrix}{\bm{J}}_{k}\bm{f}+({\bm{J}}_{\zeta}{\bm{J}}_{k}\\ +\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}\end{matrix}$
@@ VI-A Dynamical Environment with Separable State Space
Similar to (11), the following requirement should be satisfied
@@
Compared to the previous derivation (11), the drift term $\psi(\bm{q},\bm{z})$ has an additional source from the uncontrollable state’s motion $J_{z}(\bm{q},\bm{z})\dot{\bm{z}}$. In practice, we can estimate the velocity of the uncontrollable states by the finite difference or by using a state observer. A comparison of different velocity observations is shown in Section VII-B. We can derive the ATACOM-controller similarly to Eq. (12) where ${\bm{J}}_{u}$ and ${\bm{\psi}}$ should be adapted. We illustrate the effect of drift clipping with an example involving a dynamic moving obstacle in a 2D environment.
@@ VI-B Second-Order Dynamics
The previous method focuses on first-order affine control systems with state constraints, but it cannot be directly applied to higher-order systems by simply converting them to first order. We first show through an example that using only position input violates the assumptions. Then, we present a solution for second-order systems, which can be easily extended to higher-order systems.
@@
Consider an acceleration-controlled system, i.e., $\ddot{\bm{s}}={\bm{u}}_{s}$, which can be written as a first-order affine system
@@
For the constraint $k(\bm{s})<{\bm{0}}$ that only takes $\bm{s}$ as input, the Jacobian with respect to $\widehat{\bm{s}}$ is ${\bm{J}}_{k}=\begin{bmatrix}\frac{\partial}{\partial s}k&{\bm{0}}\end{bmatrix}$. Assumption 4 does not hold as $({\bm{J}}_{G})=\mathrm{rank}({\bm{J}}_{k}\bm{G})=0$. ∎
Consider a general second-order control affine system
@@
$k^{*}(\bm{s},\dot{\bm{s}})=\zeta(k(\bm{s}))+\dot{k}(\bm{s})=\zeta(k(\bm{s}))+{\bm{J}}_{k}\dot{\bm{s}}\leq{\bm{0}}$ (18)
where $\zeta(\cdot)$ is a class $\mathcal{K}$ function, following the idea of High-order Control Barrier Function (CBF) from [36]. We again introduce the slack variable $\bm{\mu}$ to construct the constraint manifold and set the time-derivatives of the new constraint $\bm{k}^{*}$ to zero and obtain
$({\bm{J}}_{\zeta}+\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}+{\bm{J}}_{k}\bm{f}+{\bm{J}}_{k}\bm{G}{\bm{u}}_{s}+\bm{A}\bm{\mu}={\bm{0}}$
By grouping the terms together, we obtain a requirement in the same form as Eq. (11). The Jacobian ${\bm{J}}_{u}$ and the drift ${\bm{\psi}}$ are defined as
${\bm{J}}_{u}=[{\bm{J}}_{k}\bm{G}\;\;\bm{A}],\;{\bm{\psi}}={\bm{J}}_{k}\bm{f}+({\bm{J}}_{\zeta}{\bm{J}}_{k}+\nabla_{s}{\bm{J}}_{k}\dot{\bm{s}})\dot{\bm{s}}$
@@ VI-C Equality Constraints
ATACOM controller can be extended for equality constraint $l(\bm{s})={\bm{0}}$, where $l:\mathcal{S}\rightarrow\mathbb{R}^{L}$ is of class $C^{1}$ with the Jacobian $J_{l}:\mathcal{S}\rightarrow\mathbb{R}^{L\times S}$. We assume the system with equality constraint is under-constrained, i.e., $L<S$. We can formulate the constraint manifold
@@
We can again construct the ATACOM controller using Eq. (12). The controlled system will approach $\mathcal{M}$ from the neighborhood $\Omega_{\eta}$ and stay on the manifold.
However, simply choosing a matrix ${\bm{T}}\in\mathbb{R}^{N\times U}$ whose diagonal entry is ${\bm{T}}_{[i,i]}=1$ is no longer a valid choice for the coordinate space in this setting. For example, for a 1-sphere in 2D, the coordinate frame ${\bm{T}}=\begin{bmatrix}1&0\end{bmatrix}^{\intercal}$ is not a valid coordinate as there are multiple points mapped to the same point. The 2-sphere does not even contain a smooth varying basis, according to the Hairy Ball Theorem [76]. Unfortunately, the choice of the reference matrix is not trivial. We leave this discussion for future work.
@@ VII Experiments
In this section, we will analyze the effectiveness of different techniques introduced in previous sections in some simple environments to provide a thorough understanding of the algorithm. In the first experiment, we will compare the effect of different slack dynamics functions and their hyperparameters in a 2D static collision avoidance environment (Fig. 7a). Then, we will compare the effects of different velocity observation methods in a dynamic environment containing moving obstacles (Fig. 7b). Additionally, we evaluate how the dynamic mismatch will impact the performance of ATACOM in a simulated air hockey task (Fig. 7d). Finally, we show in the real-world robot air hockey experiment (Fig. 7e) that we can fine-tune the RL agent from online real-world interactions with the help of ATACOM. Videos of experiments are available at https://puzeliu.github.io/TRO-ATACOM.
The experiments demonstrated in this paper majorly focus on simple tasks with low-dimensional dynamics. Nevertheless, experiments of applying ATACOM to more complex, high-dimensional, dynamic environments can be found in [13]. We demonstrate that ATACOM enables the TIAGo robot, using a unicycle model, to safely navigate to the target while avoiding blindly moving mobile robots. Furthermore, using the learned Signed Distance Function (SDF) [77] as constraints, ATACOM allows the TIAGo robot to avoid collisions with a human while holding a cup vertically in a shared workspace or to reach the target inside a shelf with complex geometry.
@@
VII-A Different Slack Dynamics in 2D-StaticEnv
We introduced multiple types of slack dynamics functions to ensure safety in Section V-A. In this experiment, we will compare the effect of different types of slack dynamics functions (Linear and Exponential) and their hyperparameters $\beta$ in a 2D collision avoidance environment, shown in Fig. 7a. A circular obstacle shown in red stays in the middle of the environment. The task is to control the planar robot (blue point) to reach the target (green point). The robot is controlled by velocity $\begin{bmatrix}\dot{x}\;\dot{y}\end{bmatrix}^{\intercal}=\begin{bmatrix}u_{x}\;u_{y}\end{bmatrix}^{\intercal}$. The constraints are collision avoidance w.r.t. the fixed obstacle, i.e., $\|{\bm{p}}_{r}-{\bm{p}}_{o}\|>r_{\text{o}}$, and keeping the robot inside the boundaries, i.e., $l_{i}<p_{r,i}<u_{i},i\in\{x,y\}$. The reward is the negative distance to the target $r=-\|{\bm{p}}_{r}-{\bm{p}}_{t}\|$. We use SAC [78] as RL algorithm. The control frequency is set to be $100Hz$ in all experiments. The learning parameters are the same among all experiments except for the parameters in comparison, i.e., the type of slack dynamics function and its parameters $\beta$. Details of the hyperparameters can be found in Table III in Appendix B. We ran 25 seeds for each experiment setting.
@@ VII-A Different Slack Dynamics in 2D-StaticEnv
The learning curves are shown in Fig. 8. In all experiment settings, the robot has not collided with the obstacle during training. The robot can reach the target in all experiment settings within a fixed horizon length except for the setting of linear slack dynamic function with $\beta=0.3$. The reason is that the action space is heavily morphed by the slack dynamics function and the robot is taking a conservative behavior. The lower reward result indicates that the robot is reaching the target slower than the higher reward result. We can clearly see that the slack dynamics function with less morphing of the action space (higher $\beta$ and exponential slack dynamics function) leads to better training performance. However, a higher $\beta$ parameter requires a higher control frequency as the manifold will have a higher curvature, causing the system to deviate from the manifold due to the discretization error.
@@
VII-B Different Velocity Observations in 2D-DynamicEnv
@@
We use a 2D Dynamic Environment as a study example, as shown in Fig. 7b. The robot (blue point) is controlled by velocity. The constraints are defined as keeping the distance to each obstacle bigger than a threshold, i.e., $\|{\bm{p}}_{r}-{\bm{p}}_{o_{i}}\|>r_{o},i=1,2,\cdots$. The target (green square) is randomly placed in the environment for each episode. We consider two types of moving obstacles: (1) The obstacles move in a fixed pattern, i.e., each obstacle moves along a circle around its initial position. (2) The obstacles move randomly in the environment. In this experiment, we use a hand-crafted policy that tries to reach the target with a linear attractor
@@
In this experiment, we compare the safety performance of (a) exact velocity observations (EXACT), (b) no zero velocity (NONE), and (c) velocities obtained from the finite difference in position (FD). In the FD setting, we added a Gaussian noise with a standard deviation of $0.03$ to the position. We evaluate the performance with different numbers of obstacles (2, 6, 10) and different obstacle speeds. In the LOW-velocity setting, the maximum speed of the obstacles is $50\%$ of the robot’s maximum speed. The MEDIUM and HIGH-velocity settings are $100\%$ and $150\%$, respectively. We ran 1000 episodes for each experiment setting. The experiment hyperparameters are in Table III in Appendix B.
@@
VII-C Safe Control with Disturbances in QuadrotorEnv
In this section, we compare ATACOM with CBF-QP in a quadrotor environment. The task is to control the Crazyflie-2 robot to track the target that moves in an $8$-shape while avoiding collisions with walls, as shown in Fig. 7c. We use the quaternion-based nominal dynamics of the quadrotor following [79]. The state space is $13$-dimensional, and the control space is the 4-dimensional. To account for environmental disturbances, we introduce a 3-dimensional translational Gaussian noise. The reward function is $r=\exp(-10\|{\bm{p}}-{\bm{p}}_{t}\|^{2})$ with ${\bm{p}}$ and ${\bm{p}}_{t}$ being the position of the quadrotor and the target. To ensure safety, the position of the quadrotor should stay with the boundary ${\bm{p}}$ $\leq{\bm{p}}\leq\bar{{\bm{p}}}$. Designing a valid constraint considering the input limit is a challenging problem. We use the following constraint function to derive a controller
[…]
$\lambda_{r}(\underline{${\bm{p}}$}_{i}-{\bm{p}}_{i})\leq k_{i}+\zeta_{r}\dot{k}_{i}\leq\lambda_{r}({\bm{p}}_{i}-{\bm{p}}_{i})\quad i\in\{0,1\}$
where the first constraint ensures the position in $z$-direction within the boundary. The second constraint restricts the attitude of the robot with respect to the distance to the boundary in $x,y$-directions. Here $k_{i}=\langle R_{z}(\bar{q}),{\bm{n}}_{i}\rangle$ is the projection of the $z$-axis in the robot frame $R_{z}(\bar{q})$ to the normal direction of the boundary ${\bm{n}}_{i}$ and $\bar{q}$ is the quaternion of the robot. $\zeta_{p}$, $\zeta_{r}$, $\lambda_{r}$ are some constant bigger than 0.
We compared CBF-QP with two optimizers, OSQP (non-differentiable solver [80] and CvxpyLayer (differentiable) [81]. The results are illustrated in Fig. 9. ATACOM demonstrates safer behavior and greater robustness to disturbances compared to the CBF method. This superiority arises because ATACOM adjusts the desired action as it approaches the boundary, whereas the QP-based method only modifies the desired action after a violation occurs. Additionally, ATACOM is computationally more efficient than CBF-QP. The gradient between the actual action and the sampled action is directly computable from an explicit mapping, while CBF-QP relies on an implicit optimization process, making gradient computation more resource-intensive. However, the heavier computation in the RL algorithm makes the speed-up effect marginal, such as saving 7 minutes from a training of 5 hours.
Figure 9: Comparison of ATACOM and CBF was conducted using different optimizers: OSQP (non-differentiable) and CvxpyLayers (differentiable). The plots in solid colors represent results obtained from the ideal simulator, while the faded plots, marked with an asterisk (*), depict results from simulations with disturbances. The Reward represents the discounted cumulative reward of the final policy. The Safe Ratio indicates the proportion of safe episodes to the total number of episodes during the entire training process. Time measures the elapsed time for each action step.
VII-D Dynamic Mismatch in Air Hockey Simulation